# Artificial barriers for stochastic differential equations and for construction of boundary-preserving schemes

### v2

@@ 3.2. Reflected SDEs
Next, we provide an overview of reflected SDEs (RSDEs), we refer to [34] for a more complete introduction to the topic of RSDEs. We define RSDEs for the Skorokhod map $\Gamma$ associated with $\operatorname{cl}D=[a,b]$ for $|a|+|b|<\infty$, and the other considered domains are treated analogously. We first provide a short intuitive description and the idea of reflected SDEs and then we provide a proper definition. In this section, instead of imposing the assumptions in Section 2, we assume that $f,g:\mathbb{R}\to\mathbb{R}$ are globally Lipschitz functions.
For simplicity, we start with the case of one barrier at 00 ($a=0$ and $b=\infty$) with continuous sample paths. The goal in this case is to define a mathematical object $\xi$ with three properties:
@@
    $\xi(t)\geq 0$ for all $t\in[0,T]$ and $t\mapsto\xi(t)$ is continuous almost surely.
@@
    $\xi$ is reflected back into the domain $(0,\infty)$ whenever $\xi$ hits 00 to avoid $\xi$ entering $(-\infty,0)$.
@@
almost surely, $L$ is non-decreasing, $L(0)=0$ and
(17) $L(t)=\int_{0}^{t}\mathbb{1}_{\xi(s)=0}\mathop{}\!\mathrm{d}L(s).$
Note that the property (17) states that the process $L(t)$ can only increase when $\xi(t)=0$ and that $L(t)$ then acts in (16) as a positive driving force to push $\xi$ into $(0,\infty)$. We use the letter $L$ to emphasise that it represent a lower barrier process. Moreover, under Lipschitz conditions on $f$ and $g$, there exists a unique solution to the Skohorod problem. We refer to [34] for more details. The considered setting in Section 2 reduces to the Lipschitz case by Assumption 2.
Next, we provide the proper mathematical definition of reflected SDEs with two barriers at $a$ and $b$ (with $|a|+|b|<\infty$).
@@
with reflection at $a$ and $b$, with $-\infty<a<b<\infty$, and the initial condition $\xi(0)=\xi_{0}\in[a,b]$ if the following holds:
@@
    $\int_{0}^{t}\mathbb{1}_{\xi(s)\in[a,b)}\mathop{}\!\mathrm{d}U(s)=\int_{0}^{t}\mathbb{1}_{\xi(s)\in(a,b]}\mathop{}\!\mathrm{d}L(s)=0$ for every $t\in[0,T]$.
@@
    $\xi(t)=\xi(0)+\int_{0}^{t}f(\xi(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\xi(s))\mathop{}\!\mathrm{d}B(s)-U(t)+L(t),\ \forall t\in[0,T]$, almost surely.
@@
    All the above integral are well-defined.
Note that if $w$ solves the Itô SDE
(18) $w(t)=\xi_{0}+\int_{0}^{t}f(\Gamma(w)(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\Gamma(w)(s))\mathop{}\!\mathrm{d}B(s),$
@@
If $\xi$ cannot hit one of the boundary points (for example $b$) then the corresponding barrier process (in this example $U$) can be completely removed or set equal to 00 (in this example $U\equiv 0$) in the definition above. This follows from uniqueness of solutions to reflected SDEs, see below or [34] for more details. Of particular interest in this paper is the case where $\xi$ cannot hit either of the boundary points (see Assumption 2) which implies that both barrier processes will, in fact, be 00 (meaning that $U\equiv 0\equiv L$), this enables us to connect the original SDE in (1) to reflected SDEs (see Proposition 21).
@@
for every $p\geq 2$, for every $t,s\in[0,T]$, where $L_{f}$ and $L_{g}$ are the Lipschitz constants of $f$ and $g$, respectively.
@@ 4.1. Construction of artificial barriers
The method of artificial barriers for constructing boundary-preserving numerical schemes is based on the idea that if the solution process $X$ to an SDE cannot leave a domain, then we can augment the SDE with barriers outside this domain to obtain a reflected SDE whose solution equals the original process $X$ almost surely. The obtained reflected SDE can then be approximated using any boundary-preserving numerical scheme for reflected SDEs. To straight away apply this idea to the considered setting would be problematic because numerical schemes for reflected SDEs are in general non-strict boundary-preserving; that is, the boundary points are not unattainable by numerical schemes for reflected SDEs. This is an issue since $f$ and $g$ that we consider typically vanish on the boundary, and most numerical schemes that hit the boundary will thereafter stay at the boundary. Instead, shortly put, we modify the coefficient functions $f$ and $g$ to move the boundary slightly inside the original domain $D$ and augment the modified SDE with artificial barriers. We can then discretise the resulting (modified) RSDE using a (non-strict) boundary-preserving numerical scheme. Note that the exact details of adding artificial barriers depend on the specific problem at hand, the here outlined method is constructed to fit our purposes. The rest of this section is dedicated to formulate the method of artificial barriers mathematically.
The first step of the method of artificial barriers is to modify the drift and diffusion coefficients such that the unattainable boundary points are moved to $a+1/N$ and $b-1/N$. If $b=\infty$, then we interpret $b-1/N=\infty$. To formalise this, we introduce an integer $N\in\mathbb{N}$ and consider the following modified Itô SDE
[…]
where $D_{N}=(a+1/N,b-1/N)\subset D$ and where $f_{N}$ and $g_{N}$ are defined below in such a way that $X_{N}(t)\in D_{N}$ for every $t\in[0,T]$ almost surely. For ease of presentation and to avoid technicalities regarding the initial value, we assume that $x_{0}\in D_{1}$. $x_{0}\in D_{1}$ can be relaxed to $x_{0}\in D$ by defining $X_{N}(0)=\Pi_{N-1}(x_{0})\in D_{N}$, where $\Pi_{N-1}:\mathbb{R}\to\operatorname{cl}D_{N-1}\subset D_{N}$ is the projection onto $\operatorname{cl}D_{N-1}$ (see (27) for a definition). Changing $x_{0}$ to $\Pi_{N-1}(x_{0})$ in (19) would only introduce a deterministic error of order $1/N$ and would not change the convergence rates.
@@
As $x\mapsto\Psi_{N}(x)$ in (20) is a linear function (both for $b=\infty$ and for $|b|<\infty$), the local behaviour of $f_{N}$ near $a+1/N$ and near $b-1/N$ is the same as local behaviour of $f$ near $a$ and near $b$, respectively. The same is true for $g_{N}$ and $g$. In particular, the boundary behaviour of $X$ (see Proposition 2) is the same as the boundary behaviour of $X_{N}$ (see Proposition 20). Observe that the case $b=\infty$ in (20) can be obtained from the case $|b|<\infty$ in (20) by
@@
The above is only well-defined if $N>\frac{2}{b-a}$. If $|b-a|>2$, then there is no issue. If $|b-a|\leq 2$, then we simply restrict $N\in\mathbb{N}$ to $N>\frac{2}{b-a}$. Henceforth, we assume, for ease of presentation, that $|b-a|>2$.
@@
If Assumption 1 is satisfied, then there exist constants $\hat{L}_{f}$ and $\hat{L}_{g}$ such that
$\left|f_{N}(x)-f_{N}(y)\right|\leq\hat{L}_{f}|x-y|$
$\left|g_{N}(x)-g_{N}(y)\right|\leq\hat{L}_{g}|x-y|$
for all $x,y\in D$ and for every $N\in\mathbb{N}$. Moreover, $\hat{L}_{f}$ and $\hat{L}_{g}$ can be chosen as
$\hat{L}_{f}=\begin{cases}L_{f},\text{ if }b=\infty,\\ L_{f}\left|\frac{b-a}{b-a-2}\right|,\text{ if }|b|<\infty,\end{cases}$
@@
$\hat{L}_{g}=\begin{cases}L_{g},\text{ if }b=\infty,\\ L_{g}\left|\frac{b-a}{b-a-2}\right|,\text{ if }|b|<\infty.\end{cases}$
@@
$\left|f_{N}(x)-f_{N}(y)\right|$ $=\left|f(\Psi_{N}(x))-f(\Psi_{N}(y))\right|\leq L_{f}\left|\Psi_{N}(x)-\psi_{N}(y)\right|$
$=L_{f}\left|\frac{b-a}{b-a-2/N}\right||x-y|$
$\leq L_{f}\left|\frac{b-a}{b-a-2}\right||x-y|$
and we let
$\hat{L}_{f}=L_{f}\left|\frac{b-a}{b-a-2}\right|.$
Alternatively, Lemma 12 follows from that the composition of two Lipschitz continuous functions is also a Lipschitz continuous function and that
(21) $\operatorname{Lip}(\Psi_{N})=\sup_{x\in\mathbb{R}}\left|\Psi_{N}^{\prime}(x)\right|=\begin{cases}1,\text{ if }b=\infty,\\ \left|\frac{b-a}{b-a-\frac{2}{N}}\right|,\text{ if }|b|<\infty,\end{cases}\leq\begin{cases}1,\text{ if }b=\infty,\\ \left|\frac{b-a}{b-a-2}\right|,\text{ if }|b|<\infty.\end{cases}$
The following lemma quantifies the error between $f$ and $f_{N}$ and $g$ and $g_{N}$, respectively, which will be used for the comparison of $X$ and $X_{N}$ in Proposition 19 below.
@@
$\left|f(x)-f_{N}(x)\right|\leq C(L_{f},a,b)N^{-1}(1+|x|)$
$\left|g(x)-g_{N}(x)\right|\leq C(L_{g},a,b)N^{-1}(1+|x|)$
for all $x\in D$ and for every $N\in\mathbb{N}$. If $b=\infty$, then the constants are independent of $a$ and $b$.
@@
(22) $\begin{split}\left|\Psi_{N}(x)-x\right|&=\left|\frac{2}{N(b-a)-2}x-\frac{b+a}{N(b-a)-2}\right|\\ &\leq N^{-1}\left(\frac{2}{|b-a-2/N|}|x|+\frac{|b+a|}{|b-a-2/N|}\right)\\ &\leq N^{-1}\left(\frac{2}{|b-a|}|x|+\frac{|b+a|}{|b-a|}\right)\leq N^{-1}C(a,b)(1+|x|)\end{split}$
@@
$\left|\Psi_{N}(x)-x\right|=\frac{1}{N}$
in the case $b=\infty$ and the constants in the lemma can be chosen to be independent of $a$ and $b$. ∎
The above definitions of $f_{N}$ and $g_{N}$ and Lemma 12 implies, in particular, that $X_{N}\in D_{N}$ is well-defined and that $X_{N}\to X$ in $L^{p}(\Omega)$ as $N\to\infty$ and we formalise this in the following proposition.
@@
${\mathbb{E}}\left[\sup_{t\in[0,T]}|X_{N}(t)|^{p}\right]\leq\begin{cases}C(p,T,L_{f},L_{g},a,x_{0}),\text{ if }b=\infty,\\ \max(|a|^{p},|b|^{p}),\text{ if }|b|<\infty,\end{cases}$
@@
Moreover, the constant $C(p,T,L_{f},L_{g},a,b,x_{0})$ is independent of $N\in\mathbb{N}$.
@@
Existence and uniqueness of a solution follows from Proposition 2 (up to modifying $f_{N}$ and $g_{N}$ outside $D_{N}$). By Lemma 12, the moment bound constant is independent of $N\in\mathbb{N}$. As $\Psi_{N}$ defined in (20) is a linear function, the local behaviour of $X_{N}$ near $\partial D_{N}$ is the same as the local behaviour of $X$ near $\partial D$. Thus,
@@
The proof of convergence $X_{N}\to X$ in $L^{p}(\Omega)$ (see Proposition 19 below) uses the two following lemmas and Grönwall’s inequality.
@@
$\left|\int_{0}^{t}f(X(s))-f_{N}(Y(s))\mathop{}\!\mathrm{d}s\right|^{p}\leq C(p,T,L_{f})\left(\int_{0}^{t}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+N^{-p}\left(1+\sup_{s\in[0,t]}|Y(s)|^{p}\right)\right).$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\int_{0}^{t}f(X(s))-f_{N}(Y(s))\mathop{}\!\mathrm{d}s\right|^{p}\right]$ $\leq C(p,T,L_{f})\int_{0}^{t}{\mathbb{E}}\left[\sup_{t\in[0,s]}|X(t)-Y(t)|^{p}\right]\mathop{}\!\mathrm{d}s$
$+C(p,T,L_{f})N^{-p}\left(1+{\mathbb{E}}\left[\sup_{t\in[0,T]}|Y(t)|^{p}\right]\right).$
@@
The estimates in Lemma 15 are also true with $\sup_{s\in[0,t]}|Y(s)|^{p}$ replaced with $\sup_{s\in[0,t]}|X(s)|^{p}$. In this case, the constant depends on the Lipschitz constant of $f_{N}$ on $D$ (which can be chosen to be independent of $N$) instead of $L_{f}$.
@@
$\leq C(p,T,L_{f})\left(\int_{0}^{t}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+N^{-p}\left(1+\sup_{s\in[0,t]}|Y(s)|^{p}\right)\right),$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\int_{0}^{t}g(X(s))-g_{N}(Y(s))\mathop{}\!\mathrm{d}s\right|^{p}\right]\\ \leq C(p,T,L_{g})\left(\int_{0}^{r}{\mathbb{E}}\left[|X(s)-Y(s)|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{s\in[0,r]}|Y(s)|^{p}\right]\right)\right).$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\int_{0}^{t}g(X(s))-g_{N}(Y(s))\mathop{}\!\mathrm{d}s\right|^{p}\right]\\ \leq C(p,T,L_{g})\left(\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}|X(t)-Y(t)|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{t\in[0,T]}|Y(t)|^{p}\right]\right)\right).$
@@
In the same way as for Lemma 15, the estimates in Lemma 17 are also true with $\sup_{s\in[0,t]}|Y(s)|^{p}$ replaced with $\sup_{s\in[0,t]}|X(s)|^{p}$. In this case, the constant depends on the Lipschitz constant of $g_{N}$ on $D$ (which can be chosen to be independent of $N$) instead of $L_{g}$.
@@
$\int_{0}^{r}\left|g(X(s))-g_{N}(Y(s))\right|^{p}\mathop{}\!\mathrm{d}s$ $\leq C(p)\left(L_{g}^{p}\int_{0}^{r}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+\int_{0}^{r}|g(Y(s))-g_{N}(Y(s))|^{p}\mathop{}\!\mathrm{d}s\right)$
$\leq C(p,L_{g})\left(\int_{0}^{r}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+N^{-p}\left(1+\sup_{s\in[0,r]}|Y(s)|^{p}\right)\right),$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\int_{0}^{t}\left|g(X(s))-g_{N}(Y(s))\right|^{p}\mathop{}\!\mathrm{d}s\right]\\ \leq C(p,T,L_{g})\left(\int_{0}^{r}{\mathbb{E}}\left[|X(s)-Y(s)|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{s\in[0,r]}|Y(s)|^{p}\right]\right)\right).$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|X_{N}(t)-X(t)\right|^{p}\right]\\ \leq C(p,T,L_{f},L_{g})\left(\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}\left|X(t)-X_{N}(t)\right|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{s\in[0,T]}|X_{N}(s)|^{p}\right]\right)\right).$

### v3

@@ 3.2. Reflected SDEs
Next, we provide an overview of reflected SDEs (RSDEs), we refer to [42] for a more complete introduction to RSDEs. We define RSDEs for the Skorokhod map $\Gamma$ associated with $\operatorname{cl}D=[a,b]$, for $|a|+|b|<\infty$, and the other considered domains are treated analogously. We first provide a short intuitive description and the idea of RSDEs, and then we provide a proper definition. In this section, instead of imposing the assumptions in Section 2, we assume that $f,g:\mathbb{R}\to\mathbb{R}$ are globally Lipschitz functions.
For simplicity, we start with the case of one barrier at 0 ($a=0$ and $b=\infty$) with continuous sample paths. The goal in this case is to define a stochastic process $\xi$ with the following three properties:
@@
    $\xi(t)\geq 0$ for all $t\in[0,T]$ and $t\mapsto\xi(t)$ is continuous, almost surely.
@@
    $\xi$ is reflected back into the domain $(0,\infty)$ whenever $\xi$ hits 0 to avoid $\xi$ entering $(-\infty,0)$.
@@
for all $t\in[0,T]$, almost surely, $L$ is non-decreasing, $L(0)=0$ and
(17) $L(t)=\int_{0}^{t}\mathbb{1}_{\xi(s)=0}\mathop{}\!\mathrm{d}L(s),\ t\in[0,T].$
Note that the property (17) states that the process $L(t)$ can only increase when $\xi(t)=0$ and that $L(t)$ then acts in (16) as a positive driving force to push $\xi$ into $(0,\infty)$. We use the letter $L$ to emphasise that it represent a lower barrier process. Moreover, under Lipschitz conditions on $f$ and $g$, there exists a unique solution to the Skorokhod problem. We refer to [42] for more details. The considered setting in Section 2 reduces to the Lipschitz case by Assumption 2.
Next, we provide the proper mathematical definition of RSDEs with two barriers at $a$ and $b$ (with $|a|+|b|<\infty$).
@@
with reflection at $a$ and $b$, with $\infty<a<b<\infty$, and the initial condition $\xi(0)=\xi_{0}\in[a,b]$ if the following holds:
@@
    $\int_{0}^{t}\mathbb{1}_{\xi(s)\in(\infty,b)}\mathop{}\!\mathrm{d}U(s)=\int_{0}^{t}\mathbb{1}_{\xi(s)\in(a,\infty)}\mathop{}\!\mathrm{d}L(s)=0$, for every $t\in[0,T]$.
@@
    $\xi(t)=\xi(0)+\int_{0}^{t}f(\xi(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\xi(s))\mathop{}\!\mathrm{d}B(s)-U(t)+L(t)$, for every $t\in[0,T]$, almost surely.
@@
    All the above integrals are well-defined.
Note that if $w$ solves the Itô functional SDE
(18) $w(t)=\xi_{0}+\int_{0}^{t}f(\Gamma(w)(s))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\Gamma(w)(s))\mathop{}\!\mathrm{d}B(s),\ t\in[0,T],$
@@
If $\xi$ cannot hit one of the boundary points (for example $b$) then the corresponding barrier process (in this example $U$) can be completely removed or set equal to 0 (in this example $U\equiv 0$) in the definition above. This follows from uniqueness of solutions to RSDEs, see below or [42] for more details. Of particular interest in this paper is the case where $\xi$ cannot hit either of the boundary points (see Assumption 2), which implies that both barrier processes will, in fact, be 0 (meaning that $U\equiv 0\equiv L$), this enables us to connect the original SDE in (1) to RSDEs (see Proposition 21).
@@
for every $p\geq 2$, for every $t,s\in[0,T]$.
@@ 4.1. Construction of artificial barriers
The method of artificial barriers for constructing boundary-preserving numerical schemes is based on the idea that if the solution process $X$ to an SDE cannot leave a domain, then we can augment the SDE with barriers outside this domain to obtain an RSDE whose solution equals the original process $X$, almost surely. The obtained RSDE can then be approximated using any boundary-preserving numerical scheme for RSDEs. To straight away apply this idea to the considered setting would be problematic because numerical schemes for RSDEs are in general non-strict boundary-preserving; that is, the boundary points are not unattainable by numerical schemes for RSDEs. This is an issue since $f$ and $g$ that we consider typically vanish on the boundary, and most numerical schemes that hit the boundary will thereafter stay at the boundary. Instead, shortly put, we modify the coefficient functions $f$ and $g$ to move the boundary slightly inside the original domain $D$ and augment the modified SDE with artificial barriers. We can then discretise the resulting (modified) RSDE using a (non-strict) boundary-preserving numerical scheme. Note that the exact details of adding artificial barriers depend on the specific problem at hand, the here outlined method is constructed to fit our purposes. The rest of this section is dedicated to formulate the method of artificial barriers mathematically.
The first step of the method of artificial barriers is to modify the drift and diffusion coefficients such that the unattainable boundary points are moved to $a+1/N$ and b1/Nb-1/N. If $b=\infty$, then we interpret $b-1/N=\infty$. To formalise this, we introduce an integer $N\in\mathbb{N}$ and consider the following modified Itô SDE
(19) $\left\{\begin{aligned} &\mathop{}\!\mathrm{d}X_{N}(t)=f_{N}(X_{N}(t))\mathop{}\!\mathrm{d}t+g_{N}(X_{N}(t))\mathop{}\!\mathrm{d}B(t),\ t\in(0,T],\\ &X_{N}(0)=x_{0}\in D_{N},\end{aligned}\right$.
where $D_{N}=(a+1/N,b-1/N)\subset D$ and where $f_{N},g_{N}:D_{N}\to\mathbb{R}$ are defined below in such a way that $X_{N}(t)\in D_{N}$, for every $t\in[0,T]$, almost surely. For ease of presentation and to avoid technicalities regarding the initial value, we assume that $x_{0}\in D_{1}$. $x_{0}\in D_{1}$ can be relaxed to $x_{0}\in D$ by defining $X_{N}(0)=\Pi_{N-1}(x_{0})\in D_{N}$, where $\Pi_{N-1}:\mathbb{R}\to\operatorname{cl}D_{N-1}\subset D_{N}$ is the projection onto $\operatorname{cl}D_{N-1}$ (see (28) for a definition). Replacing $x_{0}$ with $\Pi_{N-1}(x_{0})$ in (19) would only introduce a deterministic error of order $1/N$, and would not alter the convergence rates.
@@
As $x\mapsto\Psi_{N}(x)$ in (20) is a linear function (both for $b=\infty$ and for $|b|<\infty$), the decay of $f_{N}$ near $a+1/N$ and near b1/Nb-1/N is the same as the decay of $f$ near $a$ and near $b$, respectively. The same is true for $g_{N}$ and $g$. This implies that the boundary behaviour of $X$ (see Proposition 2) is the same as the boundary behaviour of $X_{N}$ (see Proposition 20 below). Observe that the case $b=\infty$ in (20) can be obtained from the case $|b|<\infty$ in (20) by
@@
The above is only well-defined if $N>\frac{2}{b-a}$. If $|b-a|\leq 2$, then we simply restrict $N\in\mathbb{N}$ to $N>\frac{2}{b-a}$. Henceforth, we assume, for ease of presentation, that |ba|>2|b-a|>2.
@@
If Assumption 1 is satisfied, then there exist constants $\hat{L}_{f}$ and $\hat{L}_{g}$ such that the following Lipschitz conditions hold
$\left|f_{N}(x)-f_{N}(y)\right|\leq\hat{L}_{f}|x-y|,\ \forall x,y\in D_{N},$
@@
$\left|g_{N}(x)-g_{N}(y)\right|\leq\hat{L}_{g}|x-y|,\ \forall x,y\in D_{N},$
for every $N\in\mathbb{N}$. Moreover, $\hat{L}_{f}$ and $\hat{L}_{g}$ can be chosen as
$\hat{L}_{f}=\begin{cases}\operatorname{Lip}(f),\text{ if }b=\infty,\\ \operatorname{Lip}(f)\left|\frac{b-a}{b-a-2}\right|,\text{ if }|b|<\infty,\end{cases}$
and
$\hat{L}_{g}=\begin{cases}\operatorname{Lip}(g),\text{ if }b=\infty,\\ \operatorname{Lip}(g)\left|\frac{b-a}{b-a-2}\right|,\text{ if }|b|<\infty.\end{cases}$
and such that the following linear growth conditions hold
$|f_{N}(x)|\leq\hat{L}_{f}(1+|x|),\ x\in D_{N},$
and
$|g_{N}(x)|\leq\hat{L}_{g}(1+|x|),\ x\in D_{N},$
for every $N\in\mathbb{N}$.
@@
$\left|f_{N}(x)-f_{N}(y)\right|$ $=\left|f(\Psi_{N}(x))-f(\Psi_{N}(y))\right|\leq\operatorname{Lip}(f)\left|\Psi_{N}(x)-\Psi_{N}(y)\right|$
$=\operatorname{Lip}(f)\left|\frac{b-a}{b-a-2/N}\right||x-y|$
$\leq\operatorname{Lip}(f)\left|\frac{b-a}{b-a-2}\right||x-y|.$
Alternatively, as the composition of two Lipschitz continuous functions is a Lipschitz continuous function, Lemma 12 also follows from
[…]
The following lemma quantifies the error introduced by replacing $f$ and $g$ with $f_{N}$ and $g_{N}$, respectively, which will be used for the comparison between $X$ and $X_{N}$ in Proposition 19 below.
@@
$\left|f(x)-f_{N}(x)\right|\leq C(L_{f},a,b)N^{-1}(1+|x|),\ x\in D,$
$\left|g(x)-g_{N}(x)\right|\leq C(L_{g},a,b)N^{-1}(1+|x|),\ x\in D,$
for every $N\in\mathbb{N}$. If $b=\infty$, then the constants are independent of $a$, $b$, and $x$.
@@
(22) $\begin{split}\left|x-\Psi_{N}(x)\right|&=\left|\frac{b+a}{N(b-a)-2}-\frac{2}{N(b-a)-2}x\right|\\ &\leq N^{-1}\left(\frac{|b+a|}{|b-a-2/N|}+\frac{2}{|b-a-2/N|}|x|\right)\\ &\leq N^{-1}\left(\frac{|b+a|}{|b-a-2|}+\frac{2}{|b-a-2|}|x|\right)\leq N^{-1}C(a,b)(1+|x|)\end{split}$
@@
$\left|x-\Psi_{N}(x)\right|=\frac{1}{N}$
in the case $b=\infty$ and the constants in the lemma can be chosen to be independent of $a$, $b$, and $x$. ∎
As a result of Lemma 12 and Lemma 13, we can bound the linear growth constant for $f_{N}$, independently of $N\in\mathbb{N}$:
(23) $|f_{N}(x)|\leq|f_{N}(x)-f(x)|+|f(x)|\leq C(L_{f},a,b)(1+|x|),\ x\in D.$
The same holds for $g_{N}$.
From the above lemmas implies, in particular, that the domain of (19) is $D_{N}$ and its solution $X_{N}$ enjoys moment bounds, uniform in $N\in\mathbb{N}$.
@@
${\mathbb{E}}\left[\sup_{t\in[0,T]}|X_{N}(t)|^{p}\right]\leq\begin{cases}C(p,T,L_{f},L_{g},x_{0}),\text{ if }b=\infty,\\ \max(|a|^{p},|b|^{p}),\text{ if }|b|<\infty,\end{cases}$
@@
Moreover, the constant $C(p,T,L_{f},L_{g},x_{0})$ is independent of $N\in\mathbb{N}$.
@@
Existence and uniqueness of a solution follows from Proposition 2 (up to modifying $f_{N}$ and $g_{N}$ outside $D_{N}$). By (23), the moment bound constant is independent of $N\in\mathbb{N}$. As $\Psi_{N}$ defined in (20) is a linear function, the local behaviour of $X_{N}$ near $\partial D_{N}$ is the same as the local behaviour of $X$ near $\partial D$. Thus,
@@
The proof of convergence $X_{N}\to X$ in $L^{p}(\Omega)$, as $N\to\infty$, in Proposition 19 below uses the two following lemmas and Grönwall’s inequality.
@@
$\left|\int_{0}^{t}f(X(s))-f_{N}(Y(s))\mathop{}\!\mathrm{d}s\right|^{p}\leq C(p,T,L_{f},a,b)\left(\int_{0}^{t}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+N^{-p}\left(1+\sup_{s\in[0,t]}|Y(s)|^{p}\right)\right).$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\int_{0}^{t}f(X(s))-f_{N}(Y(s))\mathop{}\!\mathrm{d}s\right|^{p}\right]$ $\leq C(p,T,L_{f},a,b)\int_{0}^{t}{\mathbb{E}}\left[\sup_{t\in[0,s]}|X(t)-Y(t)|^{p}\right]\mathop{}\!\mathrm{d}s$
$+C(p,T,L_{f},a,b)N^{-p}\left(1+{\mathbb{E}}\left[\sup_{t\in[0,T]}|Y(t)|^{p}\right]\right).$
If $b=\infty$, then the constants are independent of $a$ and $b$.
@@
The estimates in Lemma 15 are also true with $\sup_{s\in[0,t]}|Y(s)|^{p}$ replaced with $\sup_{s\in[0,t]}|X(s)|^{p}$.
@@
$\leq C(p,T,L_{f},a,b)\left(\int_{0}^{t}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+N^{-p}\left(1+\sup_{s\in[0,t]}|Y(s)|^{p}\right)\right),$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\int_{0}^{t}g(X(s))-g_{N}(Y(s))\mathop{}\!\mathrm{d}B(s)\right|^{p}\right]\\ \leq C(p,T,L_{g},a,b)\left(\int_{0}^{r}{\mathbb{E}}\left[|X(s)-Y(s)|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{s\in[0,r]}|Y(s)|^{p}\right]\right)\right).$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\int_{0}^{t}g(X(s))-g_{N}(Y(s))\mathop{}\!\mathrm{d}B(s)\right|^{p}\right]\\ \leq C(p,T,L_{g},a,b)\left(\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}|X(t)-Y(t)|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{t\in[0,T]}|Y(t)|^{p}\right]\right)\right).$
If $b=\infty$, then the constants are independent of $a$ and $b$.
@@
In the same way as for Lemma 15, the estimates in Lemma 17 are also true with $\sup_{s\in[0,t]}|Y(s)|^{p}$ replaced with $\sup_{s\in[0,t]}|X(s)|^{p}$.
@@
$\int_{0}^{r}\left|g(X(s))-g_{N}(Y(s))\right|^{p}\mathop{}\!\mathrm{d}s$ $\leq C(p,L_{g},a,b)\left(\int_{0}^{r}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+\int_{0}^{r}|g(Y(s))-g_{N}(Y(s))|^{p}\mathop{}\!\mathrm{d}s\right)$
$\leq C(p,L_{g},a,b)\left(\int_{0}^{r}|X(s)-Y(s)|^{p}\mathop{}\!\mathrm{d}s+N^{-p}\left(1+\sup_{s\in[0,r]}|Y(s)|^{p}\right)\right),$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\int_{0}^{t}\left|g(X(s))-g_{N}(Y(s))\right|^{p}\mathop{}\!\mathrm{d}s\right]\\ \leq C(p,T,L_{g},a,b)\left(\int_{0}^{r}{\mathbb{E}}\left[|X(s)-Y(s)|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{s\in[0,r]}|Y(s)|^{p}\right]\right)\right).$
Lemma 15 and Lemma 17 are the main ingredients in the following proposition.
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|X_{N}(t)-X(t)\right|^{p}\right]\\ \leq C(p,T,L_{f},L_{g},a,b)\left(\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}\left|X(t)-X_{N}(t)\right|^{p}\right]\mathop{}\!\mathrm{d}s+N^{-p}\left(1+{\mathbb{E}}\left[\sup_{s\in[0,T]}|X_{N}(s)|^{p}\right]\right)\right).$